You are the manager of a monopoly. A typical consumerโ€™s inverse demand function for your firmโ€™s product is ๐‘ƒ = 250 โˆ’ 4๐‘„ and your cost function is ๐ถ(๐‘„) = 10๐‘„. a) Determine the profit-maximizing quantity, and price under standard monopoly pricing. b) Determine the optimal two-part pricing strategy. c) Suppose you want to use block pricing strategy. How many units should you put in a block and what price you should charge for the block? d) Compare your profits from parts (a), (b) and (c)


1
Expert's answer
2021-09-23T09:11:14-0400

Solution:

a.). Profit maximizing quantity and price under monopoly is determined where MR = MC

First derive TR:

TR = P x Q

P = 250 โ€“ 4Q

TR = (250 โ€“ 4Q)"\times" Q = 250Q โ€“ 4Q2

Derive MR:

MR = "\frac{\partial TR} {\partial Q}" = 250 โ€“ 8Q

TC = 10Q

Derive MC:

MC = "\frac{\partial TC} {\partial Q}" = 10

Set MR = MC

250 โ€“ 8Q = 10

250 โ€“ 10 = 8Q

240 = 8Q

Q = 30

Substitute to get price:

P = 250 โ€“ 4Q = 250 โ€“ 4(30) = 250 โ€“ 120 = 130

P = 130

Profit = TR โ€“ TC

Profit = (130 "\times" 30) โ€“ (10 "\times" 30) = 3,900 โ€“ 300 = 3,600

Profit under standard monopoly pricing = 3,600

ย 

b.). The optimal two-part pricing strategy:

Set the price equal to marginal cost and charge a fee equal to the consumer surplus at that price.

P = MC = 10

Therefore, the per-unit price will be $10, which equals marginal cost.

Substituting this into the demand equation:

10 = 250 โ€“ 4Q

4Q = 250 โ€“ 10

4Q = 240

Q = 60

Consumer surplus = 0.5 (250 โ€“ 10) (60) = 7,200

The monopoly will charge an additional fee of 7,200 under the two-part pricing strategy.

ย 

c.). Block-pricing strategy:

TR = P x Q

P = 250 โ€“ 4Q

TR = (250 โ€“ 4Q)"\times" Q = 250Q โ€“ 4Q2

Derive MR:

MR = "\frac{\partial TR} {\partial Q}" = 250 โ€“ 8Q

TC = 10Q

Derive MC:

MC = "\frac{\partial TC} {\partial Q}" = 10

Substituting this into the demand equation:

10 = 250 โ€“ 4Q

4Q = 250 โ€“ 10

4Q = 240

Q = 60

Optimal quantity to package = 6 0

Derive the optimal price for the package:

= 0.5(250 โ€“ 10) (60) + (10 x 60) = 7,200 + 600 = 7,800

Profit under block pricing strategy = 0.5(250 โ€“ 10) (60) + (10 "\times" 60) โ€“ (10 "\times" 60) = 7,800 โ€“ 600 = 7,200

ย 

d.). The profits under the two-part pricing strategy and block pricing strategy is more profitable compared to standard monopoly pricing. Therefore, price discrimination is paramount under the two-part pricing and block pricing strategy.